# PainFormer: a Vision Foundation Model for Automatic Pain Assessment

### v2

@@ IV-A4 GSR
Regarding GSR, we observe an apparent difference in performance among the four representations. Waveform-based representations significantly outperform the others, with an initial accuracy of $87.75\%$ in the $200$-epoch configuration, surpassing other metrics by over $14\%$. During longer training sessions, there is a slight improvement across representations, suggesting that the GSR modality may have reached its performance plateau. The spectrogram-phase is the most informative among spectrograms, achieving final accuracy, recall, and F1 scores of $76.41\%$, $77.23\%$, and $76.47\%$, respectively. The waveform representation is the most effective, achieving the highest scores of $88.99\%$ for accuracy, $89.55\%$ for recall, and $88.88\%$ for the F1 metric. This variation in performance can be attributed to the inherent nature of the GSR signal. As shown in Fig. 4, GSR typically appears as a smooth curve with gradual slopes, reflecting the slow and steady changes in skin conductivity due to variations in sweat gland activity triggered by stress or arousal. In contrast, EMG signals display sharp spikes and erratic fluctuations, indicative of the rapid electrical activity associated with skeletal muscle contractions. Similarly, ECG signals are characterized by distinct cyclical patterns such as the P and T waves and the QRS complex. This suggests that the absence of more complicated patterns in GSR is not well-suited for spectral and frequency domain analysis, which is better leveraged by spectrograms. However, waveform representations yield the best results for GSR and outperform all other modalities and visual representations, underscoring their effectiveness in capturing essential physiological information from GSR signals. Table VIII summarizes the results for the GSR modality.
@@ IV-A5 Fusion
In exploring video modality fusion, four combinations were assessed: RGB and thermal, RGB and depth, thermal and depth, and the aggregation of all three—RGB, thermal, and depth. For the RGB and thermal video combination, regardless of the fusion technique employed, performance was inferior to using RGB alone, with the highest accuracy of $75.66\%$ attained via decision fusion. Similarly, the RGB and depth video combination yielded the best results with decision fusion, with a $75.53\%$ accuracy rate, yet this still fell short of the performance achieved with RGB alone. Interestingly, the fusion of thermal and depth videos marked an improvement over using depth alone, which was more effective when used independently. With decision fusion, this combination achieved a $73.02\%$ accuracy rate, surpassing depth alone by $1.35\%$. Other metrics also observed improvements, with recall and F1 scores reaching $74.46\%$ and $72.59\%$, respectively. Finally, the three video inputs of RGB thermal and depth were the only combination that led to improved results compared to the RGB in isolation. Even though some improvements were observed with the addition method, the best results came through the decision fusion, with accuracy, recall, and F1 scores of $76.55\%$, $77.91\%$, and $76.11\%$, respectively—marking improvements of $0.26\%$, $0.35\%$, and $0.55\%$. It is noted that in all experiments involving video representations, decision fusion consistently outperformed the addition method.
@@ V Comparison with existing methods
Regarding the biosignals, it is noted that in ECG-based studies, PainFormer achieved the highest results in the literature, with an accuracy of $75.49\%$ using the spectrogram-PSD representation. Compared to the subsequent leading studies [87, 46], our approach shows a significant improvement, outperforming them by more than $6\%$ and $8\%$, respectively. In the relatively few EMG-based studies employing the addition of waveform and spectrogram-PSD representations, we yielded a $72.84\%$ accuracy, significantly outperforming the following closest study [47], which achieved $63.10\%$. In the GSR-based studies, our approach, utilizing solely waveform representation and achieving an $88.99\%$ accuracy, leads to performance. It is also observed that studies using raw biosignals rather than extracting domain-specific features generally exhibit better results. The second [88] and third [89] ranked studies achieved $85.56\%$ and $84.80\%$ accuracy, respectively. The Table XII presents the corresponding results for the biosignals.
@@
Finally, concerning the AI4Pain dataset, PainFormer achieved $53.67\%$ accuracy using the RGB video modality, outperforming [91] with $49.00\%$ but falling short of [92], which reached $55.00\%$ using a transformer-based masked autoencoder. Utilizing only the fNIRS, an accuracy of $52.60\%$ was achieved, which is approximately $1\%$ lower than the $53.66\%$ reported in [93]. In a multimodal approach combining videos and waveform representations, an accuracy of $55.69\%$ was attained, surpassing [94] by more than $9\%$ and establishing the highest performance on this specific dataset to date. Table XIV presents the corresponding results.
@@
[…]
@@
[…]
@@
[…]
  [47] GSR domain-specific RF 71.90
[…]
@@
[…]
@@
Video fNIRS Fusion
@@
55.69
  • ENS: Ensemble Classifier
@@ VI Interpretation
VII Discussion

### v3

@@ IV-A4 GSR
Regarding GSR, we observe an apparent difference in performance among the four representations. Waveform-based representations significantly outperform the others, with an initial accuracy of $87.75\%$ in the $200$-epoch configuration, surpassing other metrics by over $14\%$. During longer training sessions, there is a slight improvement across representations, suggesting that the GSR modality may have reached its performance plateau. The spectrogram-phase is the most informative among spectrograms, achieving final accuracy, recall, and F1 scores of $6.41\%, 77$$7.23\%$, and $76.47\%$, respectively. The waveform representation is the most effective, achieving the highest scores of $88.99\%$ for accuracy, $89.55\%$ for recall, and $88.88\%$ for the F1 metric. This variation in performance can be attributed to the inherent nature of the GSR signal. As shown in Fig. 4, GSR typically appears as a smooth curve with gradual slopes, reflecting the slow and steady changes in skin conductivity due to variations in sweat gland activity triggered by stress or arousal. In contrast, EMG signals display sharp spikes and erratic fluctuations, indicative of the rapid electrical activity associated with skeletal muscle contractions. Similarly, ECG signals are characterized by distinct cyclical patterns such as the P and T waves and the QRS complex. This suggests that the absence of more complicated patterns in GSR is not well-suited for spectral and frequency domain analysis, which is better leveraged by spectrograms. However, waveform representations yield the best results for GSR and outperform all other modalities and visual representations, underscoring their effectiveness in capturing essential physiological information from GSR signals. Table VIII summarizes the results for the GSR modality.
@@ IV-A5 Fusion
In exploring video modality fusion, four combinations were assessed: RGB and thermal, RGB and depth, thermal and depth, and the aggregation of all three—RGB, thermal, and depth. For the RGB and thermal video combination, regardless of the fusion technique employed, performance was inferior to using RGB alone, with the highest accuracy of $75.66\%$ attained via decision fusion. Similarly, the RGB and depth video combination yielded the best results with decision fusion, with a $75.53\%$ accuracy rate, yet this still fell short of the performance achieved with RGB alone. Interestingly, the fusion of thermal and depth videos marked an improvement over using depth alone, which was more effective when used independently. With decision fusion, this combination achieved a $73.02\%$ accuracy rate, surpassing depth alone by $1.35\%$. Other metrics also observed improvements, with recall and F1 scores reaching $74.46\%$ and $72.59\%$, respectively. Finally, the three video inputs of RGB thermal and depth were the only combination that led to improved results compared to the RGB in isolation. Even though some improvements were observed with the addition method, the best results came through the decision fusion, with accuracy, recall, and F1 scores of $6.55\%, 77$$7.91\%$, and $76.11\%$, respectively—marking improvements of $0.26\%$, $0.35\%$, and $0.55\%$. It is noted that in all experiments involving video representations, decision fusion consistently outperformed the addition method.
@@ V Comparison with existing methods
Regarding the biosignals, it is noted that in ECG-based studies, PainFormer achieved the highest results in the literature, with an accuracy of $75.49\%$ using the spectrogram-PSD representation. Compared to the subsequent leading studies [92, 49], our approach shows a significant improvement, outperforming them by more than $6\%$ and $8\%$, respectively. In the relatively few EMG-based studies employing the addition of waveform and spectrogram-PSD representations, we yielded a $72.84\%$ accuracy, significantly outperforming the following closest study [41], which achieved $57.90\%$. In the GSR-based studies, our approach, utilizing solely waveform representation and achieving an $88.99\%$ accuracy, leads to performance. It is also observed that studies using raw biosignals rather than extracting domain-specific features generally exhibit better results. The second [93] and third [94] ranked studies achieved $85.56\%$ and $84.80\%$ accuracy, respectively. The Table XII presents the corresponding results for the biosignals.
@@
Finally, concerning the AI4Pain dataset, PainFormer achieved $53.67\%$ accuracy using the RGB video modality, outperforming [96] with $49.00\%$ but falling short of [97], which reached $55.00\%$ using a transformer-based masked autoencoder. Utilizing only the fNIRS, an accuracy of $52.60\%$ was achieved, which is approximately $1\%$ lower than the $53.66\%$ reported in [98]. In a multimodal approach combining videos with waveform representations, an accuracy of $55.69\%$ was achieved, surpassing [66] nearly $1\%$, [67] by more than $\%, [99$] by over $4\%$, and [100] (using only EDA) by $0.52\%$, establishing the highest performance reported on this dataset to date. Table XIV presents the corresponding results.
@@
[89] raw 2D CNN, biLSTM 69.25
[87] optical flow RF 70.20
[90] raw 2D CNN 71.00
[34] raw Vision-MLP 71.03
[101] raw 2D CNN 71.30
[41] facial landmarks, 3D distances RF 71.60
[102] facial 3D distances Deep RF 72.10
[102] facial action descriptors Deep RF 72.40
[59] facial landmarks, 3D distances RF 72.70
[50] fiducial points GNN 73.20
[51] raw Transformer 73.28
[103] raw 2D CNN, GRU 73.90
[49] raw Transformer 77.10
[91] facial landmarks STAGCN 77.30
[53] raw 3D CNN 77.50
[52] raw, rPPG 3D CNN 78.90
[53] raw, heart rate 3D CNN 88.10
Our raw Transformer 76.55
@@
[55] ECG domain-specific LR 57.40
[104] ECG domain-specific LR 57.69
[105] ECG domain-specific SVM 58.39
[106] ECG domain-specific SVM 58.62
[94] ECG raw 1D CNN, biLSTM 61.20
[41] ECG domain-specific RF 62.00
[59] ECG domain-specific SVM 62.40
[107] ECG raw 2D CNN, biLSTM 63.20
[53] ECG heart rate 3D CNN 65.00
[49] ECG heart rate Transformer 67.04
[92] ECG domain-specific FCN 69.40
Our ECG raw Transformer 75.49
@@
Our EMG raw Transformer 72.84
  [41] GSR domain-specific RF 73.80
[104] GSR domain-specific LR 74.21
[59] GSR domain-specific RF 74.40
[105] GSR domain-specific LSTM 76.86
[108] GSR domain-specific RF 80.40
[56] GSR domain-specific RF 81.90
[55] GSR domain-specific LR 82.36
[109] GSR domain-specific SVM 83.30
[107] GSR raw 1D CNN, biLSTM 83.60
[110] GSR domain-specific MLP 84.22
[54] GSR raw 1D CNN 84.57
[94] GSR raw 1D CNN, biLSTM 84.80
[93] GSR raw 1D CNN 85.56
Transformer
[111] GSR raw 1D CNN 86.21
Transformer
Our GSR raw Transformer 88.99
@@
[41] ECG, EMG, GSR domain-specific RF 74.10
[105] ECG, EMG, GSR domain-specific LSTM 77.21
[41] Video1, ECG2, EMG2, GSR2 facial landmarks1 RF 77.80
3D distances1
domain-specific2⋇
[59] Video1, ECG2, GSR2 facial landmarks1 RF 78.90
3D distances1
domain-specific2⋇
[49] Video1, ECG2 raw1, heart rate2 Transformer 82.74
[56] Video1, ECG2, EMG2, GSR2 geometric1 RF 83.10
appearance1
domain-specific2
[55] ECG, EMG, GSR domain-specific LR 83.20
[112] ECG, EMG, GSR domain-specific biLSTM 83.30
[113] ECG, EMG, GSR raw DDCAE 83.99
[114] ECG, EMG, GSR raw DDCAE, 84.25
NN
[54] ECG, EMG, GSR raw 2D CNN 84.40
[115] GSR, ECG domain-specific NN 84.58
[107] Video, GSR raw 2D CNN 84.80
biLSTM
[94] ECG, GSR raw 1D CNN 84.80
biLSTM
[57] ECG, EMG, GSR domain-specific RF 85.70
[116] ECG, EMG, GSR domain-specific RF 85.80
[58] Video1, ECG2, EMG2, GSR2 facial descriptors1 RF 86.00
domain-specific2
[95] GSR, ECG domain-specific NN 87.06
Our Video, GSR raw Transformer 89.08
@@
Video fNIRS EDA BVP Resp SpO2
@@
53.67
Our Transformer 55.69
  • BVP: Blood Volume Pulse Resp: Respiration Signal SpO2: Peripheral Oxygen Saturation ENS: Ensemble Classifier MoE: Mixture of Experts $\dagger$: AI4PAIN-First Multimodal Sensing Grand Challenge $\ddagger$: AI4PAIN-Second Multimodal Sensing Grand Challenge
@@ VI Interpretation
VII Computational Resources and Runtime
The development of Painformer and all presented experiments was conducted on a workstation running Ubuntu 22.04 LTS, with Python 3.9.18, Torch 1.11.0, and CUDA 12.9. The hardware configuration featured an NVIDIA RTX 3090 GPU paired with an AMD Ryzen 9 5950X 16-core CPU. Pre-training the foundation model using the proposed multi-task learning approach took approximately $6$ months, while pain-related tasks required 22-77 days, depending on the number of modalities and the selected fusion strategy.
VIII Discussion